Job Order Cost Accounting Entries for a Service Business

Media Connect Inc. provides advertising services for clients across the nation. Media Connect is presently working on four projects, each for a different client. Media Connect accumulates costs for each account (client) on the basis of both direct costs and allocated indirect costs. The direct costs include the charged time of professional personnel and media purchases (air time and ad space). Overhead is allocated to each project as a percentage of media purchases. The predetermined overhead rate is 60% of media purchases. On April 1, the four advertising projects had the following accumulated costs:

April 1 Balances
First Bank $61,300
Reliable Airlines 18,400
Motel 26 42,900
Blue Mountain Beverages 26,400

During April, Media Connect incurred the following direct labor and media purchase costs related to preparing advertising for each of the four accounts:

Direct Labor Media Purchases
First Bank $43,500 $160,500
Reliable Airlines 19,400 141,200
Motel 26 85,500 103,100
Blue Mountain Beverages 97,400 77,200

At the end of April, both the First Bank and Reliable Airlines campaigns were completed. The costs of completed campaigns are added to the cost of services account.

Determine each of the following for the month:

a. Direct labor costs.
$

b. Media purchases.
$

c. Overhead applied.
$

d. Cost of completed First Bank and Reliable Airlines campaigns.

Total Costs
First Bank $
Reliable Airlines $
Answer #1
a
Direct labor costs 245800 =43500+19400+85500+97400
b
Media purchases 482000 =160500+141200+103100+77200
c
Overhead applied 289200 =482000*60%
d
Total Costs
First Bank 361600 =61300+43500+160500+(160500*60%)
Reliable Airlines 263720 =18400+19400+141200+(141200*60%)